OpenStudy (anonymous):

what is the 6th term of the geometric sequence where a1= -4096 and a4= 64? answers: -1 4 1 -4

OpenStudy (anonymous):

an= ar^(n-1) a1 = -4096 a4 = 64 64 = -4096* r^3

OpenStudy (campbell_st):

well you need to find the common ratio... r a term in a geometric sequence is \[a_{n} = a_{1} \times r^{n - 1}\] so using the information provided you can find r \[64 = -4096 \times r^{4 -1}\] so \[\frac{64}{-4096} = r^3\] solve for r, then use the formula above to find the 6th term, n = 6, a = -4096 r =

OpenStudy (campbell_st):

it is